a father is now only three times as old as his son but eight years ago the father was five times as old as his son find their present ages
Answer by macston(5194)

s=son's age;f=father's age=3s
f-8yrs=5(s-8yrs) Substitute for f.
3s-8yrs=5(s-8yrs)
3s-8yrs=5s-40yrs Add 40 yrs to each side.
3s+32yrs=5s Subtract 3s from each side.
32yrs=2s Divide each side by 2.
16yrs=s ANSWER 1: The son is 16 years old.
f=3s=3(16yrs)=48 yrs ANSWER 2: The father is 48 years old.
CHECK:
f-8=5(s-8)
48-8=5(16-8)
40=5(8)
40=40